# Electrically Tunable and Enhanced Nonlinearity of Moiré Exciton-Polaritons in Transition Metal Dichalcogenide Bilayers

### v2

@@ Appendix C Momentum integration in direct and exchange interactions
$\phi^{s}_{\alpha\sigma}(\mathbf{p})=\sum_{n_{x}n_{y}}S^{s}_{\alpha,\bm{n}}\varphi_{n_{x}}(p_{x}\lambda)\varphi_{n_{y}}(p_{y}\lambda),$ (31)
@@
$-w_{\ell_{c}\tilde{\ell}_{v}}(\mathbf{q})I^{cv}_{\vec{n}_{x}}(q_{x})I^{cv}_{\vec{n}_{y}}(q_{y})-w_{\ell_{v}\tilde{\ell}_{c}}(\mathbf{q})I^{vc}_{\vec{n}_{x}}(q_{x})I^{vc}_{\vec{n}_{y}}(q_{y})\Big{]}\delta_{\ell_{v}\ell_{v}^{\prime}}\delta_{\ell_{c}^{\prime}\tilde{\ell}_{c}}\delta_{\ell_{c}\tilde{\ell}_{c}^{\prime}}\delta_{\tilde{\ell}_{v}\tilde{\ell}_{v}^{\prime}}.$ (32)
@@
$I^{\xi}_{\vec{m}}(q)=\int d^{2}k\prod_{i=1}^{4}\frac{\sqrt{\lambda_{i}}}{\sqrt{m_{i}!2^{m_{i}}\sqrt{\pi}}}\exp\Big{[}-\tfrac{1}{2}\lambda_{i}^{2}(k-u^{\xi}_{q,i})^{2}\Big{]}H_{m_{i}}(\lambda_{i}(k-u^{\xi}_{q,i})),$ (33)
where $\xi=cc,cv,vc,vv$, and $\vec{m}=[m_{1},\dots,m_{4}]$, $\vec{\lambda}=[\lambda_{1},\dots,\lambda_{4}]$ and $\vec{u}^{\xi}(q)=[u^{\xi}_{1}(q),\dots,u_{4}^{\xi}(q)]$. In Eq. (32),
$\vec{u}^{cc}(q)=\begin{bmatrix}q,&0,&q,&0\end{bmatrix}+\vec{g},\quad$ $\vec{u}^{vv}(q)=\begin{bmatrix}0,&q,&q,&0\end{bmatrix}+\vec{g},$ (34)
$\vec{u}^{cv}(q)=\begin{bmatrix}0,&0,&0,&q\end{bmatrix}+\vec{g},\quad$ $\vec{u}^{vc}(q)=\begin{bmatrix}0,&0,&q,&0\end{bmatrix}+\vec{g}$ (35)
@@
$I^{\xi}_{\vec{m}}(q)=$ $\mathrm{e}^{-\frac{1}{2}\vec{\lambda}^{2}[\bar{u}_{2}^{\xi}(q)-(\bar{u}^{\xi}_{1}(q))^{2}]}\int d^{2}k\mathrm{e}^{-\frac{1}{2}\vec{\lambda}^{2}k^{2}}\prod_{i=1}^{4}\frac{\sqrt{\lambda_{i}}}{\sqrt{m_{i}!2^{m_{i}}\sqrt{\pi}}}H_{m_{i}}(\lambda_{i}(k+\bar{u}^{\xi}_{1}(q)-u_{i}^{\xi}(q))),$ (36)
@@
$I^{\xi}_{\vec{m}}(q)=$ $\mathrm{e}^{-\frac{1}{2}\vec{\lambda}^{2}[\bar{u}_{2}^{\xi}(q)-(\bar{u}^{\xi}_{1}(q))^{2}]}\Big{[}\prod_{i=1}^{4}\sum_{r_{i}=0}^{\lfloor m_{i}/2\rfloor}\sum_{t_{i}=0}^{m_{i}-2r_{i}}[\bar{u}_{1}^{\xi}(q)-v_{i}(q)]^{t_{i}}\Big{]}M^{\vec{r},\vec{t}}_{\vec{m}}(\vec{\lambda}),$ (37)
@@
$M^{\vec{r},\vec{t}}_{\vec{m}}(\vec{\lambda})=\Big{[}\prod_{i=1}^{4}\frac{\sqrt{\lambda_{i}}}{\sqrt{m_{i}!2^{m_{i}}\sqrt{\pi}}}\frac{(-1)^{r_{i}}m_{i}!(2\lambda_{i})^{m_{i}-2r_{i}}}{r_{i}!t_{i}!(m_{i}-2r_{i}-t_{i})!}\Big{(}\frac{2}{\vec{\lambda}^{2}}\Big{)}^{(m_{i}-2r_{i}-t_{i}+1)/2}\Big{]}\frac{1}{2}\Gamma\Big{(}\frac{\sum_{i=1}^{4}(m_{i}-2r_{i}-t_{i})+1}{2}\Big{)}$ (38)
Substituting Eqs. (34)-(35) into $b$ and $c$ we have
$\bar{u}^{\xi}_{1}(q)-u_{i}^{\xi}(q)=\bar{G}_{1}-g_{i}+\begin{cases}(\bar{\lambda}^{2}_{1}+\bar{\lambda}^{2}_{3})q-(\delta_{i1}+\delta_{i3})q,&\text{for }\xi=cc\\ (\bar{\lambda}^{2}_{2}+\bar{\lambda}^{2}_{3})q-(\delta_{i2}+\delta_{i3})q,&\text{for }\xi=vv\\ \bar{\lambda}^{2}_{4}q-\delta_{i4}q,&\text{for }\xi=cv\\ \bar{\lambda}^{2}_{3}q-\delta_{i3}q,&\text{for }\xi=vc\\ \end{cases}$ (39)
@@
$\bar{u}^{\xi}_{2}(q)-[\bar{u}^{\xi}_{1}(q)]^{2}=$ $\bar{G}_{2}-\bar{G}_{1}^{2}+\begin{cases}2[\bar{\lambda}_{1}^{2}(g_{1}-\bar{G}_{1})+\bar{\lambda}_{3}^{2}(g_{3}-\bar{G}_{1})]q+[\bar{\lambda}_{1}^{2}+\bar{\lambda}_{3}^{2}-(\bar{\lambda}_{1}^{2}+\bar{\lambda}_{3}^{2})^{2}]q^{2},&\text{for }\xi=cc\\ 2[\bar{\lambda}_{2}^{2}(g_{2}-\bar{G}_{1})+\bar{\lambda}_{3}^{2}(g_{3}-\bar{G}_{1})]q+[\bar{\lambda}_{2}^{2}+\bar{\lambda}_{3}^{2}-(\bar{\lambda}_{2}^{2}+\bar{\lambda}_{3}^{2})^{2}]q^{2},&\text{for }\xi=vv\\ 2\bar{\lambda}_{4}^{2}(g_{4}-\bar{G}_{1})q+[\bar{\lambda}_{4}^{2}-(\bar{\lambda}_{4}^{2})^{2}]q^{2},&\text{for }\xi=cv\\ 2\bar{\lambda}_{3}^{2}(g_{3}-\bar{G}_{1})q+[\bar{\lambda}_{3}^{2}-(\bar{\lambda}_{3}^{2})^{2}]q^{2},&\text{for }\xi=vc\end{cases}$ (40)
@@
$\Big{[}\Theta^{cc}_{\vec{t}_{x},\vec{t}_{y}}(\vec{\lambda},\vec{g})+\Theta^{vv}_{\vec{t}_{x},\vec{t}_{y}}(\vec{\lambda},\vec{g})-\Theta^{cv}_{\vec{t}_{x},\vec{t}_{y}}(\vec{\lambda},\vec{g})-\Theta^{vc}_{\vec{t}_{x},\vec{t}_{y}}(\vec{\lambda},\vec{g})\Big{]}M^{\vec{r}_{x}\vec{t}_{x}}_{\vec{n}_{x}}(\vec{\lambda})M^{\vec{r}_{y}\vec{t}_{y}}_{\vec{n}_{y}}(\vec{\lambda})\delta_{\ell_{v}\ell_{v}^{\prime}}\delta_{\ell_{c}^{\prime}\tilde{\ell}_{c}}\delta_{\ell_{c}\tilde{\ell}_{c}^{\prime}}\delta_{\tilde{\ell}_{v}\tilde{\ell}_{v}^{\prime}}.$ (41)
@@
Appendix D MoS2 homobilayer with $\theta\approx 0$
In this section, we calculate the spectrum of the MoS2 homobilayer with $\theta\approx 0$ in Fig.4. In contrast to the $\theta\approx 60^{\circ}$ case, the attractive interaction in this stacking is stronger because the inversion symmetry forbid large electric dipole moment in a hybridized exciton.
Figure 4: (a) Moiré exciton spectrum with $\theta\approx 0$. (b) the nonlinearity of moiré exciton.
Appendix E Calculation of polariton density of state $\rho_{c}(\omega)$
@@
$\bar{H}_{\mathrm{xp}}=\begin{bmatrix}\omega_{c}-i\kappa&U\\ U^{T}&H_{X}\end{bmatrix},\quad$ (43)

### v3

@@ III Momentum integration in direct and exchange interactions
$\phi^{s}_{\alpha\sigma}(\mathbf{p})=\sum_{n_{x}n_{y}}S^{s}_{\alpha,\bm{n}}\varphi_{n_{x}}(p_{x}\lambda)\varphi_{n_{y}}(p_{y}\lambda),$ (S20)
@@
$-w_{\ell_{c}\tilde{\ell}_{v}}(\mathbf{q})I^{cv}_{\vec{n}_{x}}(q_{x})I^{cv}_{\vec{n}_{y}}(q_{y})-w_{\ell_{v}\tilde{\ell}_{c}}(\mathbf{q})I^{vc}_{\vec{n}_{x}}(q_{x})I^{vc}_{\vec{n}_{y}}(q_{y})\Big{]}\delta_{\ell_{v}\ell_{v}^{\prime}}\delta_{\ell_{c}^{\prime}\tilde{\ell}_{c}}\delta_{\ell_{c}\tilde{\ell}_{c}^{\prime}}\delta_{\tilde{\ell}_{v}\tilde{\ell}_{v}^{\prime}}.$ (S21)
@@
$I^{\xi}_{\vec{m}}(q)=\int d^{2}k\prod_{i=1}^{4}\frac{\sqrt{\lambda_{i}}}{\sqrt{m_{i}!2^{m_{i}}\sqrt{\pi}}}\exp\Big{[}-\tfrac{1}{2}\lambda_{i}^{2}(k-u^{\xi}_{q,i})^{2}\Big{]}H_{m_{i}}(\lambda_{i}(k-u^{\xi}_{q,i})),$ (S22)
where $\xi=cc,cv,vc,vv$, and $\vec{m}=[m_{1},\dots,m_{4}]$, $\vec{\lambda}=[\lambda_{1},\dots,\lambda_{4}]$ and $\vec{u}^{\xi}(q)=[u^{\xi}_{1}(q),\dots,u_{4}^{\xi}(q)]$. In Eq. (S21),
$\vec{u}^{cc}(q)=\begin{bmatrix}q,&0,&q,&0\end{bmatrix}+\vec{g},\quad$ $\vec{u}^{vv}(q)=\begin{bmatrix}0,&q,&q,&0\end{bmatrix}+\vec{g},$ (S23)
$\vec{u}^{cv}(q)=\begin{bmatrix}0,&0,&0,&q\end{bmatrix}+\vec{g},\quad$ $\vec{u}^{vc}(q)=\begin{bmatrix}0,&0,&q,&0\end{bmatrix}+\vec{g}$ (S24)
@@
$I^{\xi}_{\vec{m}}(q)=$ $\mathrm{e}^{-\frac{1}{2}\vec{\lambda}^{2}[\bar{u}_{2}^{\xi}(q)-(\bar{u}^{\xi}_{1}(q))^{2}]}\int d^{2}k\mathrm{e}^{-\frac{1}{2}\vec{\lambda}^{2}k^{2}}\prod_{i=1}^{4}\frac{\sqrt{\lambda_{i}}}{\sqrt{m_{i}!2^{m_{i}}\sqrt{\pi}}}H_{m_{i}}(\lambda_{i}(k+\bar{u}^{\xi}_{1}(q)-u_{i}^{\xi}(q))),$ (S25)
@@
$I^{\xi}_{\vec{m}}(q)=$ $\mathrm{e}^{-\frac{1}{2}\vec{\lambda}^{2}[\bar{u}_{2}^{\xi}(q)-(\bar{u}^{\xi}_{1}(q))^{2}]}\Big{[}\prod_{i=1}^{4}\sum_{r_{i}=0}^{\lfloor m_{i}/2\rfloor}\sum_{t_{i}=0}^{m_{i}-2r_{i}}[\bar{u}_{1}^{\xi}(q)-v_{i}(q)]^{t_{i}}\Big{]}M^{\vec{r},\vec{t}}_{\vec{m}}(\vec{\lambda}),$ (S26)
@@
$M^{\vec{r},\vec{t}}_{\vec{m}}(\vec{\lambda})=\Big{[}\prod_{i=1}^{4}\frac{\sqrt{\lambda_{i}}}{\sqrt{m_{i}!2^{m_{i}}\sqrt{\pi}}}\frac{(-1)^{r_{i}}m_{i}!(2\lambda_{i})^{m_{i}-2r_{i}}}{r_{i}!t_{i}!(m_{i}-2r_{i}-t_{i})!}\Big{(}\frac{2}{\vec{\lambda}^{2}}\Big{)}^{(m_{i}-2r_{i}-t_{i}+1)/2}\Big{]}\frac{1}{2}\Gamma\Big{(}\frac{\sum_{i=1}^{4}(m_{i}-2r_{i}-t_{i})+1}{2}\Big{)}$ (S27)
Substituting Eqs. (S23)-(S24) into $b$ and $c$ we have
$\bar{u}^{\xi}_{1}(q)-u_{i}^{\xi}(q)=\bar{G}_{1}-g_{i}+\begin{cases}(\bar{\lambda}^{2}_{1}+\bar{\lambda}^{2}_{3})q-(\delta_{i1}+\delta_{i3})q,&\text{for }\xi=cc\\ (\bar{\lambda}^{2}_{2}+\bar{\lambda}^{2}_{3})q-(\delta_{i2}+\delta_{i3})q,&\text{for }\xi=vv\\ \bar{\lambda}^{2}_{4}q-\delta_{i4}q,&\text{for }\xi=cv\\ \bar{\lambda}^{2}_{3}q-\delta_{i3}q,&\text{for }\xi=vc\\ \end{cases}$ (S28)
@@
$\bar{u}^{\xi}_{2}(q)-[\bar{u}^{\xi}_{1}(q)]^{2}=$ $\bar{G}_{2}-\bar{G}_{1}^{2}+\begin{cases}2[\bar{\lambda}_{1}^{2}(g_{1}-\bar{G}_{1})+\bar{\lambda}_{3}^{2}(g_{3}-\bar{G}_{1})]q+[\bar{\lambda}_{1}^{2}+\bar{\lambda}_{3}^{2}-(\bar{\lambda}_{1}^{2}+\bar{\lambda}_{3}^{2})^{2}]q^{2},&\text{for }\xi=cc\\ 2[\bar{\lambda}_{2}^{2}(g_{2}-\bar{G}_{1})+\bar{\lambda}_{3}^{2}(g_{3}-\bar{G}_{1})]q+[\bar{\lambda}_{2}^{2}+\bar{\lambda}_{3}^{2}-(\bar{\lambda}_{2}^{2}+\bar{\lambda}_{3}^{2})^{2}]q^{2},&\text{for }\xi=vv\\ 2\bar{\lambda}_{4}^{2}(g_{4}-\bar{G}_{1})q+[\bar{\lambda}_{4}^{2}-(\bar{\lambda}_{4}^{2})^{2}]q^{2},&\text{for }\xi=cv\\ 2\bar{\lambda}_{3}^{2}(g_{3}-\bar{G}_{1})q+[\bar{\lambda}_{3}^{2}-(\bar{\lambda}_{3}^{2})^{2}]q^{2},&\text{for }\xi=vc\end{cases}$ (S29)
@@
$\Big{[}\Theta^{cc}_{\vec{t}_{x},\vec{t}_{y}}(\vec{\lambda},\vec{g})+\Theta^{vv}_{\vec{t}_{x},\vec{t}_{y}}(\vec{\lambda},\vec{g})-\Theta^{cv}_{\vec{t}_{x},\vec{t}_{y}}(\vec{\lambda},\vec{g})-\Theta^{vc}_{\vec{t}_{x},\vec{t}_{y}}(\vec{\lambda},\vec{g})\Big{]}M^{\vec{r}_{x}\vec{t}_{x}}_{\vec{n}_{x}}(\vec{\lambda})M^{\vec{r}_{y}\vec{t}_{y}}_{\vec{n}_{y}}(\vec{\lambda})\delta_{\ell_{v}\ell_{v}^{\prime}}\delta_{\ell_{c}^{\prime}\tilde{\ell}_{c}}\delta_{\ell_{c}\tilde{\ell}_{c}^{\prime}}\delta_{\tilde{\ell}_{v}\tilde{\ell}_{v}^{\prime}}.$ (S30)
@@
IV MoS2 homobilayer with $\theta\approx 0$
In this section, we calculate the spectrum of the MoS2 homobilayer with $\theta\approx 0$ in Fig.S1. In contrast to the $\theta\approx 60^{\circ}$ case, the attractive interaction in this stacking is stronger because the inversion symmetry forbid large electric dipole moment in a hybridized exciton.
Figure S1: (a) Moiré exciton spectrum with $\theta\approx 0$. (b) the nonlinearity of moiré exciton.
V Calculation of polariton density of state $\rho_{c}(\omega)$
@@
$\bar{H}_{\mathrm{xp}}=\begin{bmatrix}\omega_{c}-i\kappa&U\\ U^{T}&H_{X}\end{bmatrix},\quad$ (S32)
@@
VI Nonlinear absorption spectrum
The absorption for a exciton with energy $\mathcal{E}$ can be obtained by using the Fermi golden rule with a Lorentz broadening ($\beta=5$ meV), leading to
$\Gamma(\mathcal{E})=\frac{2\pi}{\hbar}\sum_{\bar{\alpha}}\frac{(\beta/\pi)|\langle 0|\mathcal{X}^{\bar{\alpha}}_{\sigma}\mathcal{H}_{\mathrm{sc}}c^{\dagger}_{\mathbf{Q}}|0\rangle|^{2}}{[\mathcal{E}^{\bar{\alpha}}_{\sigma}+n_{\bar{\alpha}}g^{\bar{\alpha}}_{X}-\mathcal{E}(\mathbf{Q})]^{2}+\beta^{2}},$ (S33)
where the photon dispersion is $\mathcal{E}(\mathbf{Q})=\hbar|\mathbf{Q}|$. Also, $n_{\bar{\alpha}}=\langle\hat{x}^{\dagger}_{\bar{\alpha}\sigma}\hat{x}_{\bar{\alpha}\sigma}\rangle/L^{2}$is the excitons density in the background which depends on pump power. This result is used in the plot of Fig.3.